OpenStudy (anonymous):

0.002 = 0.001/(1-F^2) how do u rearrange this for F? to find a final answer for F.

OpenStudy (unklerhaukus):

you could start by multiplying both sids of the equation by (1-F^2) and diving both sides by 0.002

OpenStudy (anonymous):

0.002*(1-F^2) = 0.001

OpenStudy (unklerhaukus):

right ,keep going ...

OpenStudy (anonymous):

1-F^2 = 0.001/0.002

OpenStudy (anonymous):

now minus 1 from both sides?

OpenStudy (unklerhaukus):

yeah you can do that ,

OpenStudy (anonymous):

ok so that leaves -F^2 on one side, what next unk?

OpenStudy (unklerhaukus):

well next , multiply both sides by negative 1

OpenStudy (unklerhaukus):

and simplify the right hand side

OpenStudy (anonymous):

F^2 = - (0.001/0.002) + 1 is that correct?

OpenStudy (unklerhaukus):

yep, now that fraction will simplify

OpenStudy (anonymous):

thanks unkle you are too good, thanks alot mate, i struggle with rearranging equations. i have to practice more, i got the answer in my text book. thanks mate

OpenStudy (unklerhaukus):

did you get both solutions for F?

OpenStudy (anonymous):

i got 0.707 sound right?

OpenStudy (anonymous):

0.71 rounded up

OpenStudy (unklerhaukus):

yeah thats right if you want your anser in decimal, but i would probably leave the solution as \[F=\pm\tfrac1{\sqrt2}\]

OpenStudy (anonymous):

ok, yeah this is for hydraulics so i need the answer in decimal, it is a calcs for Froudes number, do u know engineering, u would be good at it.

OpenStudy (unklerhaukus):

ah, in that case the negative solution dosent make much sense so it is fine to leave your anser as 0.707 or 0.71 i study physics so i do know a little bit about engineering , but i hadn't heard of Froudes number until now